OpenStudy (anonymous):

anyone a pro with differential equations? It has been too long and I need some help

OpenStudy (anonymous):

I feel ok with them, not pro but hopefully I can help

OpenStudy (anonymous):

\[y'''+y=0 \] where y(o)=0 , y'(0) =1, y''(0)= 0 thank you please help

OpenStudy (anonymous):

sorry I cannot help, havent done 3rd order ones.

OpenStudy (anonymous):

Ok how about find all solutions to y'-2y=1 thank you for trying

OpenStudy (anonymous):

you can multiply with an integrating factor, do you remember that?

OpenStudy (anonymous):

I will get a pen and paper

OpenStudy (anonymous):

thank you I know its alot of work

OpenStudy (anonymous):

ok I solved it

OpenStudy (anonymous):

so do you know the technique of integrating factor?

OpenStudy (anonymous):

you r the best lets see if I get it :)

OpenStudy (anonymous):

you want to make a full derivative out of the left hand side

OpenStudy (anonymous):

for this you need to have something like dy/dx+ some y=whatever

OpenStudy (anonymous):

my computer is slowing down sorry ...yes I remeber that

OpenStudy (anonymous):

no problem, what is the integrating factor here?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

after we integrate x dx on the other side

OpenStudy (anonymous):

it is \[e ^{\int\limits_{}^{} the multiplier of y}dx\]

OpenStudy (anonymous):

that is -2 here

OpenStudy (anonymous):

the integral of -2 is -2x (you dont need the constant here) so multiply through by \[e ^{-2}\]

OpenStudy (anonymous):

-2x

OpenStudy (anonymous):

Im sorry the computer wont let me type

OpenStudy (anonymous):

It keeps stopping but I am here reading...Yes I remeber now the -2x

OpenStudy (anonymous):

the result is \[e ^{-2x}\] dy/dx -2 e^{-2x} y=e ^{-2x}

OpenStudy (anonymous):

do you recognise the full derivative now?

OpenStudy (anonymous):

yes it is coming back to me

OpenStudy (anonymous):

what about when we have a trig function like y''+4y=cos x

OpenStudy (anonymous):

this is harder but my favourite :)

OpenStudy (anonymous):

better your fav than mine :) :) i have not done these in 17 years it is so hard to remember you r a big help

OpenStudy (anonymous):

first you have to think of the homogeneous equation (meaning y''+4y=0) you have to find the general solution for that, for this we use the auxiliary equation. that just says that ay''+by'+cy=0 and write a quadratic equation for it,\[a \lambda ^{2}+b \lambda +c=0\]

OpenStudy (anonymous):

where lambda=e^x

OpenStudy (anonymous):

or in the form of e^x

OpenStudy (anonymous):

here you will have to solve y^2+4=0

OpenStudy (anonymous):

do you know how to?

OpenStudy (anonymous):

I will be back in 5 min

OpenStudy (anonymous):

i have no idea Im trying to follow along with you and my book but so confused I need to watch a tutorial i think and i will be bac

OpenStudy (anonymous):

you can only solve this with complex numbers

OpenStudy (anonymous):

thank ufor your help

OpenStudy (anonymous):

you had enough? How is it that after 17 years you are doing this again?

OpenStudy (anonymous):

http://www.analyzemath.com/calculus/Differential_Equations/second_order.html This page explains it quite well

OpenStudy (anonymous):

i am a teacher but needed 2 credits of advanced math i truly didnt want to take up your time

OpenStudy (anonymous):

I will have an exam about this in 2weeks so it is not a waste of time for me

OpenStudy (anonymous):

I need alot of explanation and my computer is not cooperating ...to be honest this work is due tomorrow and my brain is blank we can continue if u want to help me..

OpenStudy (anonymous):

well it is up to you, I am free now for an hour

OpenStudy (anonymous):

Consider the equation y' + (cos x) y= e^-sinx find the solution that satisfies \[\phi\](\[\pi\] ) = pi

OpenStudy (anonymous):

what is the end bit? that isnt clear

OpenStudy (anonymous):

I guess you have to use the integrating factor here as well. it will be e^integral(cosx) so e^sinx.

OpenStudy (anonymous):

that way you get \[(e ^{sinx}y)'=1\]

OpenStudy (anonymous):

so the LHS= x+C y=(x+C)/e^sinx